OpenStudy (anonymous):

A brick falls from the top of a 5m building.Calculate the velocity with which the brick reaches the ground ? How long it takes the brik to reach the ground ?

OpenStudy (anonymous):

you have to solve this problem by conservation of energy..loss of potential energy=gain in kinetic energy. \[mgh=1/2mv^2\] here m cancels out,g=9.8,h=5 mt,v=velocity gained. you will get the velocity t which it reaches the ground.now by v=U+gt,where u=0,you will get the time t,taken to reach the ground

OpenStudy (anonymous):

if you take g=10,the answer will come easily,then v will be 10 m/s. and the time taken will be 1 sec.

OpenStudy (theeric):

@gjecii If you don't know about the kinetic and potential energy ideas, and you're studying projectile motion, you might prefer the formula\[v_f^2=v_i^2+2\ a\ d\]Here, \(v_f\) is the final velocity, \(v_i\) is the initial velocity (which is \(0\)), \(a\) is the acceleration which is acceleration due to gravity, and \(d\) is the displacement (which is the 5 meter drop)!

OpenStudy (theeric):

For the equation I wrote, you would need to solve for \(v_f\) still. So it'd be\[v_f=\sqrt{v_i^2+2\ a\ d}\] And @sagnik_das has a great response if you are aware of those ideas. The potential energy at the top of the fall is going to be the same as the kinetic energy at the bottom of the fall. So you can set them equal to each other and solve.